OpenStudy (anonymous):

Given: y=4x^3 - ln(3x+2) Find: y'(1)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@satellite73

OpenStudy (raden):

do you know what's the derivative of 4x^3 = ... also the derivative of ln(3x+2) = ...

OpenStudy (anonymous):

y'=12x^3 - 1/(3x+2) ?

OpenStudy (anonymous):

-3/(3x+2)?

OpenStudy (raden):

12x^3 or 12x^2 ? the rest is okay :)

OpenStudy (anonymous):

12x^2

OpenStudy (anonymous):

so the answer is 57/5?

OpenStudy (raden):

so, you have y'=12x^2 - 3/(3x+2) put x=1 to y' y'(x=1) = 12(1)^2 - 3/(3*1+2) = 12 - 3/5 = 57/5 you are right ;)

OpenStudy (anonymous):

but the answer at the back was 11 2/5

OpenStudy (raden):

that's same :) 11 2/5 is just mix fraction of ... :P

OpenStudy (anonymous):

ok, can you help me with another please? ^^

OpenStudy (raden):

sure if i can :)